OpenStudy (anonymous):

A rectangular storage container with an open top is to have a volume of 60 cubic meters. The length of its base is twice the width. Material for the base costs 60 dollars per square meter. Material for the sides costs 9 dollars per square meter. Find the cost of materials for the cheapest such container. This is what I have so far but the answer never seems to come out right. l=2w V=lwh=2w^2h h=60/2w^2=30/w^2 Total Surface Area=2w^2+2(2w)(30/w^2)+2w(30/w^2) =2w^2+(120/w)+(60/w) =2w^2+(180/w) Then you add in the prices of the base and sides =(60)2w^2+(9)(180/w)

OpenStudy (anonymous):

=120w^2+(1620/w) Then take the derivative =240w-(1620/w^2) Set that equal to 0 and solve for w 240w=1620/w^2 w^3=1620/240 w w=(cbrt(6.75) Then you plug that back into your equation =120(cbrt(6.75)^2)-(1620/(cbrt(6.75))) but apparently this is wrong...help please? Where did I go wrong?